# Large Language Models as Span Annotators

### v1

@@ 4 Experiments
Task # Cat. # Texts Avg. Len Task Setup Novel Data
D2T-Eval 6 1296 118/715 Kasner & Dušek (2024)
MT-Eval 2 2,854 26/185 Kocmi et al. (2024a) ×
Propaganda 18 100 914/4,659 Da San Martino et al. (2019) ×
Table 1: Overview of span annotation tasks used in our experiments. # Cat. denotes the number of categories used in the task (see Appendix E for their listings), # Texts the number of texts annotated, and Avg. Len the average number of words/characters in the output.
We cover three span annotation tasks of different quantitative and qualitative aspects: evaluation of data-to-text generation (D2T-Eval; Section 4.1), evaluation of machine translation (MT-Eval; Section 4.2), and propaganda detection (Propaganda; Section 4.3). In Table 1, we provide a quantitative overview of our tasks. [^5] See also Figure 1 and Appendix I for the visualization of examples of our tasks. Additionally, we discuss how we prevent the issue of data contamination for our tasks in Appendix C. We further describe our process of collecting annotations in Section 4.4.
@@ 4.1 Task: Evaluation of Data-to-text Generation
In D2T-Eval, we evaluate semantic accuracy and stylistic aspects of texts generated from structured data Celikyilmaz et al. (2020); Sharma et al. (2022). We apply span annotation to the text $Y$ generated from the input structured data $X$.
@@
We select three subtasks of D2T generation: generating five-day weather forecasts (openweather), generating product descriptions (gsmarena), and generating reports of soccer games (football). We use the approach of Kasner & Dušek (2024) to download novel structured data from publicly available APIs, to make sure they are unseen for all LLMs (cf. Appendix C on data contamination). We provide more details on the data collection process in Section D.1.
@@
We generate outputs for the structured inputs using LLMs. Specifically, we select two state-of-the-art models, Llama 3.3 70B Grattafiori et al. (2024) and GPT-4o Hurst et al. (2024), and two smaller models, Gemma 2 2B Team et al. (2024) and Phi 3.5 3.8B Abdin et al. (2024). We prompt the models in a zero-shot setting, asking them to generate a summary of the given data using approximately five sentences. [^6] See Sections A.2 and A.3 for details on the models and Appendix F for the prompts.
@@
  • $\mathcal{D}_{\texttt{test}}$ (1200 outputs) – the subset for LLM evaluation, annotated with crowdworkers.
  • $\mathcal{D}_{\texttt{dev}}$ (84 outputs) – the subset for the study of prompt variants, annotated internally.
@@
We gather span annotations for $\mathcal{D}_{\texttt{test}}$ with crowdworkers from Prolific.com. We apply best practices for gathering human annotations, including an iterative process of refining annotation guidelines and pre-selecting best-performing annotators using a qualification task Tseng et al. (2020); Iskender et al. (2020); Huang et al. (2023); Zhang et al. (2023). Of the 230 annotators that participated in the qualification task, we selected 50 best annotators to annotate our data. We provide more details on the process of collecting human annotations in Section D.2.
4.2 Task: Evaluation of Machine Translation
For MT-Eval, we use the dataset of system outputs from the WMT 2024 general shared task (Kocmi et al., 2024a). The system outputs were annotated with the Error Span Annotation (ESA) protocol (Kocmi et al., 2024b) using professional translators. We focus on character-level span annotations of major and minor translation errors provided in the dataset. The inputs $X$ are the source texts used to generate the translation $Y$ [^7] Note the annotations in MT-Eval cannot overlap and need not to be aligned with word boundaries.
@@ 4.2 Task: Evaluation of Machine Translation
The original dataset has nearly 50k model outputs, making it too costly to run the full evaluation campaign. Therefore, we use all available system outputs for 10 sampled inputs for each language, giving us 2,854 outputs which we use for the experiments.
@@ 4.3 Task: Propaganda Technique Detection
For the Propaganda task, we use the dataset of Da San Martino et al. (2019). The dataset contains news collected from mostly propagandistic on-line sources. The token-level annotations in the dataset were created by expert annotators and cover 18 categories of logical fallacies and persuasion techniques. We use the test split for our experiments. Note that there are no inputs $X$ for the task, as all the annotated categories are intrinsic to the evaluated text $Y$.
@@ 4.4 Collecting LLM annotations
For our span annotator LLMs, we use a mix of open and proprietary state-of-the-art models: (1) instruction-tuned models Llama 3.3 70B Grattafiori et al. (2024), GPT-4o Hurst et al. (2024), Claude 3.7 Sonnet Anthropic (2025), and (2) reasoning models: DeepSeek-R1 70B DeepSeek-AI et al. (2025), o3-mini OpenAI (2025) and Gemini 2.0 Flash Thinking Deepmind (2025).
We run the open models (Llama 3.3, DeepSeek-R1) locally in 4-bit quantization using ollama.
@@
We define several prompt variants for our experiments. $\mathcal{P}_{\texttt{base}}$ is the base prompt that includes the guidelines $\mathcal{G}$ as given to human annotators. By extending $\mathcal{P}_{\texttt{base}}$, we implement few-shot ($\mathcal{P}_{\texttt{5shot}}$) and chain-of-thought ($\mathcal{P}_{\texttt{cot}}$) prompts. We also ablate $\mathcal{P}_{\texttt{base}}$ by removing extended guidelines ($\mathcal{P}_{\texttt{noguide}}$) and not asking for explanations ($\mathcal{P}_{\texttt{noreason}}$). The full prompts can be found in Appendix F.
@@
We use factgenie Kasner et al. (2024), a tool which supports both collecting span annotations from humans via a web interface and from LLMs via API calls.
@@ 5 Results
In this section, we describe the results of our experiments on span annotation with LLMs. We evaluated the LLM annotations both automatically using the metrics from Section 3.3 (Section 5.2) and manually (Section 5.3).
@@ 5.1 Prompting Techniques
To answer RQ1, we study the differences between prompting techniques on the $\mathcal{D}_{\texttt{dev}}$ split using the open models (Llama 3.3 and DeepSeek-R1). The results are shown in Table 2.
@@
$\mathcal{P}_{\texttt{base}}$ 0.16 0.13 2.4 0.23 0.20 1.0
$\mathcal{P}_{\texttt{noguide}}$ 0.09 0.08 3.4 0.19 0.16 1.6
$\mathcal{P}_{\texttt{noreason}}$ 0.19 0.13 2.2 0.22 0.18 1.1
$\mathcal{P}_{\texttt{cot}}$ 0.08 0.10 0.8 0.21 0.19 1.1
$\mathcal{P}_{\texttt{5shot}}$ 0.23 0.18 2.5 0.20 0.16 1.4
Table 2: Comparison of different prompting techniques on the $\mathcal{D}_{\texttt{dev}}$ split with Llama 3.3 and DeepSeek-R1: $F_{1}$ score, $\gamma$ score, and the average number of annotations per output (#a/o).
Including detailed guidelines seems beneficial: omitting the guidelines ($\mathcal{P}_{\texttt{noguide}}$) lowers the performance of both models. In contrast, not letting the model explain the annotation ($\mathcal{P}_{\texttt{noreason}}$) does not have a substantial effect on the model performance. For Llama 3.3, the chain-of-thought (CoT) prompting ($\mathcal{P}_{\texttt{cot}}$) makes it produce less annotations per example than the base variant (0.76 vs 2.42), leading to lower F1 and $\gamma$ scores. [^8] The model with $\mathcal{P}_{\texttt{cot}}$ often “overthinks” the annotations, deciding not to annotate cases of errors against which it can find some arguments. We provide an example of this behavior in Table 27. Few-shot prompting ($\mathcal{P}_{\texttt{5shot}}$) brings ambivalent results, increasing Llama 3.3 scores but doing the opposite for DeepSeek-R1.999This observation is aligned with DeepSeek-AI et al. (2025), who note that few-shot prompting consistently degrades the performance of DeepSeek-R1. Following this preliminary study, we decided to focus on prompting with $\mathcal{P}_{\texttt{base}}$ for our main experiments.
@@
Figure 2: Comparison of LLMs and human annotators using $\mathcal{P}_{\texttt{base}}$. The stripe pattern denotes proprietary models, darker shades denote reasoning models. For $F_{1}$ score, the grey bar denotes its soft variant. The dashed horizontal lines are the average value of agreement between human annotators. Detailed results are included in Tables 8, 16 and 20.
To answer RQ2, we compare LLM annotations to human annotations using metrics described in Section 3.3. Wherever possible, we compare model results with an average IAA of human annotators. [^10] For D2T-Eval and MT-Eval, we estimate the bounds from the examples annotated by pairs of annotators. For Propaganda, we use the $\gamma$ score reported by Da San Martino et al. (2019) for agreement of the annotators before consolidation ($\gamma=0.31$). The overall results for all the tasks are outlined in Figure 2. We provide detailed results for individual tasks in Appendix G.
Figure 3: Confusion matrix of category counts averaged across models for D2T-Eval (Contradictory, Not checkable, Misleading, Incoherent, Repetitive, Other, cf. Table 5).
@@
Llama 3.3 0.27 0.20 0.08 0.08 0.10 0.05
GPT-4o 0.27 0.19 0.07 0.06 0.16 0.12
Claude 3.7 0.36 0.25 0.13 0.14 0.27 0.19
DeepS. R1 0.30 0.28 0.09 0.16 0.16 0.12
o3-mini 0.38 0.33 0.11 0.19 0.29 0.26
Gem. 2-FT 0.35 0.31 0.13 0.16 0.22 0.16
@@
Performance is task-dependent.
LLMs provide annotations that are aligned with human annotations to a degree, with exact agreement depending on the task. For D2T-Eval, all models surpass the average $F_{1}$-scores (both soft and hard) between human annotators (0.13 and 0.21, respectively), although only o3-mini surpasses the $\gamma$ score (0.25). Performance is notably lower for Propaganda, where all models remain below the $\gamma$ agreement of human annotators. It should be noted that expert annotators were used in this task, which can make it more difficult for the models to reach their agreement level. This task also has the largest number of categories, manifested by a large increase in the soft $F_{1}$-score as opposed to its hard variant. On average, the models are also below human performance for MT-Eval, with Claude 3.7 showing the best results (0.22 $F_{1}$-score vs. 0.25 of human annotators).
@@
DeepSeek-R1, which is a reasoning model, outperforms the same-sized instruction-tuned Llama 3.3. The superiority of DeepSeek-R1 is most pronounced on D2T-Eval ($F_{1}$-score of 0.19 vs. 0.15, $\gamma$ score of 0.19 vs. 0.11). We hypothesize that this is because statements in texts generated from structured data often involve numerical reasoning. The same observation applies to OpenAI models, where o3-mini outperforms GPT-4o despite o3-mini’s price per token being approximately 2x lower. Both findings demonstrate the potential of test-time scaling Snell et al. (2024); Welleck et al. (2024). A notable exception from this trend is Claude 3.7 Sonnet, which is mostly on par with o3-mini (although its price per token is approximately 3x higher).
Models mostly confuse related categories.
The results of category classification on D2T-Eval suggest that the models mostly tend to confuse annotation categories such as Contradictory, Not checkable and Misleading that are all related to semantic accuracy (cf. Figure 3 and Section G.8). This suggests that categorization errors may be less serious and may be related to category ambiguity or subjective understanding of category definitions. The results also show that the models use a less diverse distribution of categories, resorting to the Contradictory category more often than human annotators.
LLMs are more cost-efficient than human annotators.
Using LLMs, as opposed to human annotators, is notably different once the financial aspect is taken into account: For D2T-Eval, crowdsourced annotation for 1k outputs costs approximately \$500, while annotating the same amount of outputs with the high-performance model o3-mini LLM costs \$3.60. We provide a more detailed cost- and time-wise analysis in Section A.1.
@@ 5.3 Manual Analysis of LLM Annotations
To answer RQ3, we manually analyzed the quality of LLM annotations on 216 samples. For each model, we sampled 3 annotations per error category in D2T-Eval and 1 annotation per error category in Propaganda. We show the results in Figure 4.
Overall, 49.5% of LLM-generated annotations and 50.5% of reasons were marked as correct for both tasks (with 9.2% of annotations and 12.5% of reasons additionally marked as partially correct). Reasoning models perform better, with 56.4% of their annotations and 58.3% marked as completely correct. The annotations made by Gemini 2.0 Flash Thinking and DeepSeek-R1 were the most accurate in D2T-Eval. o3-mini performed well across both tasks, even though Propaganda proved to be challenging for all models.
@@
Qualitatively, we find that the models often select wrong error categories despite identifying real issues (e.g. labeling contradictory statements as "incoherent"). Models also tend to be overly attentive, flagging noise in the data (quotes or non-article content) as errors, or marking slight numerical variations (such as rounded values) as misleading. All of these cases could be tackled by more descriptive guidelines or additional examples. However, in some cases, the models also misread or misinterpret data (e.g., claiming wind speed measurements do not exist when they do), which hints at deeper issues with understanding the data. Sources of incorrect reasons vary from incomplete explanations (addressing only part of a multi-issue span), irrelevant explanations (e.g., counting omissions as sources of errors), to incorrectly flagging subjective statements (e.g., “long-lasting usage”) as factual errors. Occasionally, the model admits that it marked a correct span as an error, such as in “The description of the game’s duration aligns with the data, providing coherent information”. [^11] Note that this typically happens to GPT-4o. Since OpenAI API ensures JSON key ordering, the reason had to be generated before the annotation (same as for the other models).
@@
Interestingly, the LLM annotations that were marked as correct have only 24% hard character-level overlap (51% soft) with human annotations. This fact led us to analyze the quality of human annotations in D2T-Eval. On a sample of 108 human annotations, we found that only 45.3% of annotations and 54.6% of reasons were marked as correct, which is comparable to LLM annotations. This suggests that the task is hard even for human annotators, and the quality of annotations from crowdworkers varies, even if they are preselected using a qualification task.
Figure 4: Results of our manual analysis. We analyzed 18 annotations and their explanations for each model and task (216 annotations in total). The color bars show annotations that we classified as correct, partially correct, assigned wrong category, incorrect, and undecidable cases. Detailed results are provided in Tables 22 and 23 in Appendix G.

### v2

@@ 4 Experiments
Task # Cat. # Texts Avg. Len Novel Data
D2T-Eval 6 1,296 118/715
MT-Eval 2 2,854 26/185
Propaganda 18 100 914/4,659
Table 1: Overview of span annotation tasks used in our experiments. # Cat. denotes the number of categories used in the task (see Appendix C for their listings), # Texts the number of texts annotated, and Avg. Len the average number of words/characters in the output.
We cover three span annotation tasks of different quantitative and qualitative aspects: evaluating data-to-text generation (D2T-Eval; Section 4.1), identifying errors in machine translation (MT-Eval; Section 4.2), and detecting propaganda techniques (Propaganda; Section 4.3). In Table 1, we provide a quantitative overview of our tasks. [^6] See also Figure 1 and Appendix G for the visualization of examples of our tasks.
@@ 4.1 Task: Evaluation of Data-to-text Generation
In D2T-Eval, we focus on evaluating the semantic accuracy and stylistic aspects of texts generated from structured data Celikyilmaz et al. (2020); Sharma et al. (2022). The input is the text $Y$ generated from the structured data $X$.
@@
We select three D2T generation subtasks: generating five-day weather forecasts (openweather), generating product descriptions (gsmarena), and generating reports of soccer games (football). We use the approach of Kasner & Dušek (2024) to download novel structured data from publicly available APIs to make sure that they are unseen for all LLMs (cf. Section 4.4 on data contamination). We provide more details on the data collection process in Section B.1.
@@
We generate outputs for the structured inputs using LLMs. Specifically, we selected two state-of-the-art models, Llama 3.3 70B Grattafiori et al. (2024) and GPT-4o Hurst et al. (2024), and two smaller models, Gemma 2 2B Team et al. (2024) and Phi 3.5 3.8B Abdin et al. (2024). We prompt the models in a zero-shot setting, asking them to generate a summary of the given data using approximately five sentences. [^7] See Sections A.1 and A.2 for details on the models and Appendix D for the prompts.
@@
  • $\mathcal{D}_{\texttt{test}}$ (1200 outputs) – the subset for LLM evaluation, annotated with crowdworkers,
  • $\mathcal{D}_{\texttt{dev}}$ (84 outputs) – the subset for the study of prompt variants, annotated internally,
@@
We gather span annotations for $\mathcal{D}_{\texttt{test}}$ with crowdworkers from Prolific.com. We apply best practices for gathering human annotations, including an iterative process of refining annotation guidelines and preselecting best-performing annotators using a qualification task Tseng et al. (2020); Iskender et al. (2020); Huang et al. (2023); Zhang et al. (2023). Of the 230 annotators that participated in the qualification task, we selected 50 best annotators to annotate our data. We provide more details on the process of collecting human annotations in Section B.2.
4.2 Task: Identifying Errors in Machine Translation
For MT-Eval, we use the dataset of system outputs from the WMT 2024 general shared task (Kocmi et al., 2024b). The system outputs were annotated with the Error Span Annotation (ESA) protocol (Kocmi et al., 2024c) by professional translators. We focus on character-level span annotations of major and minor translation errors provided in the dataset. The inputs $X$ in MT-Eval are the source texts used to generate the translation $Y$ [^8] Note unlike for the other tasks, the annotations in MT-Eval cannot overlap and need not to be aligned with word boundaries.
@@ 4.2 Task: Identifying Errors in Machine Translation
The original dataset has nearly 50k model outputs, making it too costly to run the full evaluation campaign. Therefore, we used all available system outputs for 10 sampled inputs for each language, giving us 2,854 outputs that we use for the experiments.
@@ 4.3 Task: Propaganda Technique Detection
For the Propaganda task, we use the dataset of Da San Martino et al. (2019). The dataset contains news collected from mostly propagandistic on-line sources. The token-level annotations in the dataset created by expert annotators cover 18 categories of logical fallacies and persuasion techniques. We use the test split for our experiments. Note that there are no inputs $X$ for the task, as all the annotated categories are intrinsic to the evaluated text $Y$.
4.4 Mitigating Data Contamination
We took preventive measures to ensure that our results are not affected by data contamination Balloccu et al. (2024); Dong et al. (2024); Jiang et al. (2024b):
  • For D2T-Eval, we follow the approach of Kasner & Dušek (2024) to download a new set of structured data from publicly available sources and generate new outputs.
  • For Propaganda, the dataset is publicly available since 2019, but due to its format cannot be easily memorized. [^9] The model would need to separately match the annotations for numerical token ids with the text.
@@ 4.5 Collecting LLM annotations
For our span annotator LLMs, we use a mix of open and proprietary state-of-the-art models: (1) instruction-tuned models Llama 3.3 70B Grattafiori et al. (2024), GPT-4o Hurst et al. (2024), Claude 3.7 Sonnet Anthropic (2025), and (2) reasoning models: DeepSeek-R1 70B DeepSeek-AI et al. (2025), o3-mini OpenAI (2025) and Gemini 2.0 Flash Thinking Deepmind (2025). We run the open models (Llama 3.3, DeepSeek-R1) locally in 4-bit quantization using ollama.
@@
We define several prompt variants for our experiments. $\mathcal{P}_{\texttt{base}}$ is the base prompt that includes the guidelines $\mathcal{G}$ as given to human annotators. By extending $\mathcal{P}_{\texttt{base}}$, we implement few-shot ($\mathcal{P}_{\texttt{5shot}}$) and chain-of-thought ($\mathcal{P}_{\texttt{cot}}$) prompts. We also ablate $\mathcal{P}_{\texttt{base}}$ by removing extended guidelines ($\mathcal{P}_{\texttt{noguide}}$) and not asking for explanations ($\mathcal{P}_{\texttt{noreason}}$). The full prompts can be found in Appendix D.
@@
We use factgenie Kasner et al. (2024), a tool that supports collecting span annotations from humans via a web interface and from LLMs via API calls.
@@ 5 Results
In this section, we describe the results of our experiments on span annotation with LLMs. We first investigate the effect of prompting techniques in Section 5.1. Next, we evaluate the LLM annotations using automatic metrics (Section 5.2) and manually analyze the errors in the model outputs (Section 5.3).
@@ 5.1 Prompting Techniques
We perform a small-scale study on $\mathcal{D}_{\texttt{dev}}$ using open models (Llama 3.3 and DeepSeek-R1) to study the differences between prompting techniques. The results are shown in Table 2.
@@
$\mathcal{P}_{\texttt{base}}$ 0.20 0.13 2.4 0.25 0.20 1.0
$\mathcal{P}_{\texttt{noguide}}$ 0.11 0.08 3.4 0.20 0.16 1.6
$\mathcal{P}_{\texttt{noreason}}$ 0.22 0.13 2.2 0.24 0.18 1.1
$\mathcal{P}_{\texttt{cot}}$ 0.09 0.10 0.8 0.24 0.19 1.1
$\mathcal{P}_{\texttt{5shot}}$ 0.25 0.18 2.5 0.21 0.16 1.4
Table 2: Comparison of prompting techniques on the $\mathcal{D}_{\texttt{dev}}$ (#a/o is the average number of annotations per output).
Including detailed guidelines seems beneficial: omitting the guidelines ($\mathcal{P}_{\texttt{noguide}}$) lowers the performance of both models. In contrast, not letting the model explain the annotation ($\mathcal{P}_{\texttt{noreason}}$) does not have a substantial effect. For Llama 3.3, the chain-of-thought (CoT) prompting ($\mathcal{P}_{\texttt{cot}}$) makes it produce less annotations per example than the base variant (0.8 vs. 2.4), leading to lower F1 and $\gamma$ scores. [^10] The model with $\mathcal{P}_{\texttt{cot}}$ tends to “overthink” the annotations, deciding not to annotate cases of errors against which it can find some arguments. We provide an example of this behavior in Table 27. Few-shot prompting ($\mathcal{P}_{\texttt{5shot}}$) brings ambivalent results, increasing Llama 3.3 scores but doing the opposite for DeepSeek-R1.111111This observation is aligned with DeepSeek-AI et al. (2025), who note that few-shot prompting consistently degrades the performance of DeepSeek-R1. Given these considerations, we decided to use $\mathcal{P}_{\texttt{base}}$ for our main experiments.
@@
Figure 2: Comparison of LLMs using $\mathcal{P}_{\texttt{base}}$ and human annotators. The stripe pattern denotes proprietary models, darker shades denote reasoning models. For $F_{1}$ score, the grey bar denotes its soft variant. The dashed horizontal lines for D2T-Eval denote agreement between our human annotators. Detailed results are included in Tables 8, 16 and 20.
Next, we compute IAA between LLM and human annotations using the metrics described in Section 3.3. The overall results for all tasks are given in Figure 2. We also provide detailed results for individual tasks in Appendix E.
Figure 3: Confusion matrix for D2T-Eval (Contradictory, Not checkable, Misleading, Incoherent, Repetitive, Other, cf. Table 5).
@@
Llama 3.3 0.29 0.20 0.11 0.08 0.12 0.05
GPT-4o 0.30 0.19 0.09 0.06 0.18 0.12
Claude 3.7 0.38 0.25 0.16 0.14 0.27 0.28
DeepS. R1 0.34 0.28 0.12 0.16 0.18 0.12
o3-mini 0.42 0.33 0.15 0.19 0.29 0.31
Gem. 2-FT 0.38 0.31 0.16 0.16 0.22 0.24
@@
DeepSeek-R1 generally outperforms the same-sized Llama 3.3. The superiority of DeepSeek-R1 is most pronounced on D2T-Eval ($F_{1}$-score of 0.23 vs. 0.18, $\gamma$ score of 0.19 vs. 0.11). The same observation applies to OpenAI models, where o3-mini outperforms GPT-4o despite o3-mini’s price per token being approximately 2x lower. A notable exception to this trend is the Claude 3.7 Sonnet, which is mostly on par with o3-mini and excels on MT-Eval. However, its price per token is approximately 3x higher (cf. Table 4 for model costs).
LLMs reach human IAA on D2T-Eval, Propaganda is harder.
For D2T-Eval, we can compare model results with an average IAA on examples annotated by two human annotators (cf. Section B.2). Here, o3-mini, Claude 3.7 and Gemini 2.0 mostly reach or surpass human agreement. For Propaganda, the upper bound of IAA is the result of Da San Martino et al. (2019), who report $\gamma=0.31$ for annotators before consolidation. This score is substantially higher (the best LLM score being $\gamma=0.16$ for the Claude 3.7 Sonnet). However, this task used expert annotators and has the largest number of categories. The latter manifests itself in the large difference between the soft and hard $F_{1}$ scores. [^12] The WMT24 dataset for MT-Eval contains examples annotated with a pair of annotators, but these examples form only a small fraction and exhibit high variance between language pairs. Therefore, we decided to omit the comparison with human IAA in MT-Eval.
Models confuse related categories.
Confusion matrices (cf. Figure 3 and Section E.8) suggest that the models tend mainly to confuse related categories such as Contradictory, Not checkable and Misleading for D2T-Eval, which may be related to ambiguity or subjective understanding of category definitions. The models also use a less diverse distribution of categories, e.g., resorting to the Contradictory category more often than human annotators.
LLMs are more efficient than human annotators.
For D2T-Eval, crowdsourced annotation for 1k outputs costs approximately \$500, while annotating the same amount of outputs with the high-performance model o3-mini LLM costs \$3.60 (cf. Table 4). The crowdworkers take 129.1 seconds per output on average, which is better than DeepSeek-R1 on our local infrastructure, but order of magnitude slower than API-based models. [^13] Also note that unlike LLMs, the crowdworkers are not required to produce explanations for the annotations. This suggests that LLMs can be an effective alternative to human annotators both in terms of costs and time.
Metric crowdw. Llama 3.3 DeepSeek-R1 Claude 3.7 Sonnet o3-mini
Cost per 1k outputs (\$) 500 - - 10.5 3.6
Time per output (s) 129.1 21.6 227.5 9.0 21.8
Table 4: Estimate of costs and time requirements on D2T-Eval: crowdworkers on Prolific, open models (Llama 3.3, DeepSeek-R1), and proprietary models (Claude 3.7 Sonnet, o3-mini). Note that we omit the infrastructure and energy costs for the models deployed locally.
@@ 5.3 Manual Analysis of LLM Annotations
We manually analyzed the quality of LLM annotations on 216 samples from D2T-Eval and Propaganda. For each model, we sampled three annotations per category in D2T-Eval and one annotation per category in Propaganda. We show the results in Figure 4.
In total, we marked 49.5% of LLM-generated annotations and 50.5% of reasons as correct (with 9.2% of annotations and 12.5% of reasons additionally marked as partially correct). Reasoning models perform better, with 56.4% of their annotations and 58.3% marked as completely correct. The most accurate annotations on D2T-Eval were those made by Gemini 2.0 and DeepSeek-R1. o3-mini performed well on both tasks, although Propaganda proved to be challenging for all models.
@@
Qualitatively, we find that the models often select wrong error categories despite identifying real issues (e.g. labeling contradictory statements as "incoherent"). Models also tend to be overly attentive, flagging noise in the data (quotes or non-article content) as errors, or marking slight numerical variations (such as rounded values) as misleading. All of these cases could be tackled by more descriptive guidelines or additional examples. However, in some cases, the models also misread or misinterpreted the data (e.g., claiming wind speed measurements do not exist when they do), which hints at deeper issues with understanding the data. Incorrect explanations vary from incomplete explanations (addressing only part of a multi-issue span), irrelevant explanations (e.g., counting omissions as sources of errors) to incorrectly flagging subjective statements (e.g., “long-lasting usage”) as factual errors. Occasionally, the model admits that it marked a correct span as an error, such as in “The description of the game’s duration aligns with the data, providing coherent information”. [^14] This typically happens to GPT-4o, even though OpenAI API ensures JSON key ordering so the explanation should have been generated before the annotation (cf. Section 3.2).
@@
Interestingly, the LLM annotations that were marked as correct have only 24% hard character-level overlap (51% soft) with human annotations. This fact led us to analyze the quality of human annotations in D2T-Eval. On a sample of 108 human annotations, we marked 45.3% of the annotations and 54.6% of the explanations correct, which is comparable to the LLM annotations. This suggests that the task is hard even for human annotators, and the quality of annotations from crowdworkers varies, even if they are preselected using a qualification task.
Figure 4: Results of our manual analysis. We analyzed 18 annotations and their explanations for each model and task (216 annotations in total). The color bars show annotations that we classified as correct, partially correct, assigned wrong category, incorrect, and undecidable cases. Detailed results are provided in Tables 22 and 23 in Appendix E.